OpenStudy (seb.cal):

Please check my work :) My Question: How many moles of sulfuric acid are needed to produce 57.8 milliliters of water? Water has the density of 0.987 g/mL. Show all steps of your calculation as well as the final answer. NaOH + H2SO4 → Na2SO4 + H2O My Answer: Balanced Equation: 2NaOH + H2SO4 yieelds Na2SO4 + 2H2O. The Math: 57.8ml times 0.987 g/ml = 57.05 g H2O. 57.05g H2O over 18 g/mole = 3.17 moles H2O. 3.17moles H2O times (1H2SO4 over 2H2O) = 1.58 moles H2SO4. @bbypanda16

OpenStudy (anonymous):

57.8ml x 0.987g/ml = 57.05g H2O 57.05g H2O / 18g/mole = 3.17moles H2O 2NaOH + H2SO4 → Na2SO4 + 2H2O.....balanced 3.17moles H2O x (1H2SO4 / 2H2O) = 1.58moles H2SO4

OpenStudy (anonymous):

what do i need to do with NaOH + H2SO4 → Na2SO4 + H2O

OpenStudy (seb.cal):

@LegendaryClasher NaOH + H2SO4 → Na2SO4 + H2O is the equation they gave me to work with.. In My Answer, I had to balance that equation and use what the product was to get my final answer in moles.

OpenStudy (anonymous):

okk good you have it right

OpenStudy (seb.cal):

Cool, tysm :) @LegendaryClasher

OpenStudy (anonymous):

np